OpenStudy (anonymous):

7a-{2a-{3a+2(a-2)]}+5

OpenStudy (anonymous):

On problems like these you just start with the inner-most set of parentheses and work your way out. so 2(a-2)=2a-4 2a-(3a+2a-4)=2a-3a-2a+4=-3a+4 7a-{-3a+4+5}=7a+3a-4-5=10a-9 so your answer is 10a-9 you see how I got there?

OpenStudy (anonymous):

The above is correct until the final step. The 5 is not included in the parenthesis so it should read:\[7a - (-3a + 4) + 5\] which becomes:\[7a + 3a - 4 + 5\] which equals the answer:\[10a + 1\]

OpenStudy (anonymous):

I saw that it was outside the parentheses however there was an unclosed parentheses earlier so I assumed (possibly incorrectly) that there was supposed to be a parentheses at the end as that would make the most sense. Sorry for the mistake. :)